[figure: 9.7 cm]

THE GROUND STATE FOR THE ISING MODEL ±J ON THE

TUBE

Marian Błaszyk Institute of Physics, A. Mickiewicz University, Poznań, Poland

The Ising model with exchange integrals ±J for a square lattice spanned over infinite length tubes of the width of 2, 3 or 4 bonds is considered. The paper gives a short description of a method of linear equations allowing a reduction of a physical problem of determination of e.g. energy, entropy, magnetisation for systems with frozen disorder, to a solution of a system of linear equations. The results for energy, entropy and magnetisation possible for a short presentation are given.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : M. Błaszyk

Address for correspondence : Institute of Physics, A. Mickiewicz University, Poznań, ul. Umultowska 85, Poland

Email address : blaszyk@main.amu.edu.pl